Thread Rating:
  • 0 Vote(s) - 0 Average
  • 1
  • 2
  • 3
  • 4
  • 5
It's 4%, AKA the Go Fuck Yourself Post.
#17
0.04 is the chance of being chased, 1-0.04 is the chance of not being chased. The sum doesn't reach 1 because the numbers provided don't cover all cases.

Supposing a person is followed once, we end up with (0.04^(1)) + (1-0.04^(1)) = 1. Pretty standard stuff there.

The odds of an event occuring multiple times is (probability)^(times occuring) * (probability)^(times not occuring). The sum of separate odds is the odds of one of the summed odds occuring. Therefore, the odds of being followed twice when three targes are chosen are:
0.04^2 * 0.96 + 0.96 * 0.04^2 = 0.003072

Being followed once is:
0.96^2 + 0.04 + 0.96 * 0.04 * 0.96 + 0.04 * 0.96^2 = 0.110592

Being followed all three times is:
0.04^3 = 0.000064

Being followed 0 times is:
0.96^3 = 0.884736

Being followed 2 times, but not in a row is:
0.04 * 0.96 * 0.04 = 0.001536

And the sum of those is : 1

I was wrong the first time, hope this clarifies. Also, the odds of not being chased an entire fight that has six chases is 0.96^6 = 0.782757789696, or greater than 75%, which suits the bads ratio pretty well Big Grin.
Achievements are like the Md after someone's name.
Q: Know what they call a doctor that graduated 800th out of 800 in their class?
A: Doctor.
Reply


Messages In This Thread

Forum Jump:


Users browsing this thread: 2 Guest(s)